# ProbTalk3D: Non-Deterministic Emotion Controllable Speech-Driven 3D Facial Animation Synthesis Using VQ-VAE

### v2

@@ 3.1. Dataset: 3DMEAD
We choose 3DMEAD for our experiments as it offers a relatively large-scale, high-quality facial animation data with coverage of diverse emotions. Motion data is sampled at $25fps$. Each frame in the dataset is represented using FLAME (Li et al., 2017b) 3D Morphable Model (3DMM) parameters $\{\beta,\theta,\psi\}\in\mathbb{R}^{406}$, where $\beta\in\mathbb{R}^{300}$ denotes the face shape, $\theta_{jaw}\in\mathbb{R}^{3}$ denotes the eular angle rotation (x,y,z) of the jaw bone, $\theta_{global}\in\mathbb{R}^{3}$ denotes the global head pose, and $\psi\in\mathbb{R}^{100}$ denotes the expression parameters. However, similar to EMOTE, we utilize only $\{\psi,\theta_{jaw}\}\in\mathbb{R}^{53}$ for model training, where $\psi\in\mathbb{R}^{50}$ representing the first 50 of the total 100 expression parameters. The original training configuration of EMOTE splits the dataset into training-validation-test sets keeping all sequences per subject while having different subjects in each set. In this way, there is no ground truth data to perform a quantitative evaluation for which EMOTE only conducted a perceptual user study. In contrast, our proposed split keeps a small number of sequences from each training subject for validation and testing allowing comparison of generated samples with respect to the ground truth. Although our training is done on fewer sequences than EMOTE, given the large scale of 3DMEAD dataset, we demonstrate that this split provides sufficient information for effective training and generation of animations perceptually superior in comparison to the EMOTE model. More details about the dataset split can be found in the supplementary material.
@@ 3.2. Problem Formulation
Figure 3. Stage 2: We keep the motion autoencoder trained in the previous stage frozen and train the HuBERT based Audio Encoder in such a way that given audio, $a$, it produces audio latent, $z^{\text{a}}$ that closely resembles the motion latent, $z^{\text{m}}$ produced by the frozen motion encoder. The Style Vector, $\mathcal{C}$ (i.e. concatenated one-hot vectors containing information about subject ID, emotion class and emotion intensity) is used to learn the Style Embedding that is fused with the encoded audio information. Unlike stage 1 training, the audio latent $z^{\text{a}}$ is used to get the quantized codebook latent $z^{\prime}$ that is then decoded into facial animation $\hat{\mathcal{X}}$, utilizing the frozen motion decoder.
[…]
ProbTalk3D stage 2 detailed figure.
@@ 3.3.2. Stage 2: Speech and Emotion Conditioned
After learning the motion prior in stage 1, speech and emotion conditioned stage 2 is trained which consists of an audio encoder that encodes the continuous raw audio/speech data into discrete hidden representations and fuses the speaking style (i.e. style embedding obtained using Style Vector, $\mathcal{C}$) with the encoded audio information. As shown in Fig.3, the Motion Autoencoder trained in stage 1 is kept frozen in this stage to train the Audio Encoder in a supervised manner so that the latent representation, $z^{a}$ from the audio encoder closely resembles the latent representation, $z^{m}$ from the frozen Motion Encoder.
@@
We employ pretrained HuBERT (Hsu et al., 2021), a transformer based speech recognition model in our audio encoder, and made appropriate modifications to apply it to our downstream task of facial animation synthesis following (Haque and Yumak, 2023; Stan et al., 2023). After the last hidden layer of HuBERT, inspired by the aforementioned previous works, we adjust the input representation so it temporally aligns with the paired animation data and add a linear projection layer that projects the encoded hidden state into a 256-dimensional hidden state to be element-wisely multiplied with the Style Embedding. Style Embedding is obtained using a linear projection of concatenated one-hot vectors of the style annotations (subject ID, emotion, intensity). In our case, the style vector, $\mathcal{C}$ is a 43-dimensional vector (as we have 32 subjects, 8 basic emotions, and 3 emotion intensity categories) that is linearly transformed into a 256-dimensional Style Embedding to facilitate the fusion of audio and style information. After fusing the audio information and style information, we introduce a 1D convolutional layer followed by 12 transformer layers to get the latent representation $z^{a}$ that goes through the frozen motion decoder to obtain the predicted motion, $\hat{\mathcal{X}}$.
@@
The goal in stage 2 training is to get audio latent $z^{a}$ such that it is as close as possible to the motion latent, $z^{m}$. Therefore, we construct $\mathcal{L}_{lat}=\mathcal{L}_{1}(z^{m},z^{a})$ that computes the $\mathcal{L}_{1}$ loss between the latent representations. Additionally, we further add the reconstruction losses. The overall loss function is then defined in Eq.(5). The loss weights set for training are $\lambda_{\text{lat}}=1$, $\lambda_{\text{rec}}^{\text{exp}}=0.15$, and $\lambda_{\text{rec}}^{\text{jaw}}=0.1$.
@@ 4.1. Quantitative Evaluation
Figure 6. Comparison using heatmap visualization of mean and standard deviation of generated animation by different models together with ground truth (GT) given audio sequence, uttering the sentence: “The revolution now underway in materials handling makes this much easier”.

### v3

@@ 3.1. Dataset: 3DMEAD
We choose 3DMEAD for our experiments as it offers a relatively large-scale, high-quality facial animation data with coverage of diverse emotions. Motion data is sampled at $25fps$. Each frame in the dataset is represented using FLAME (Li et al., 2017b) 3D Morphable Model (3DMM) parameters $\{\beta,\theta,\psi\}\in\mathbb{R}^{406}$, where $\beta\in\mathbb{R}^{300}$ denotes the face shape, $\theta_{jaw}\in\mathbb{R}^{3}$ denotes the Euler angle rotation (x,y,z) of the jaw bone, $\theta_{global}\in\mathbb{R}^{3}$ denotes the global head pose, and $\psi\in\mathbb{R}^{100}$ denotes the expression parameters. However, similar to EMOTE, we utilize only $\{\psi,\theta_{jaw}\}\in\mathbb{R}^{53}$ for model training, where $\psi\in\mathbb{R}^{50}$ representing the first 50 of the total 100 expression parameters. The original training configuration of EMOTE splits the dataset into training-validation-test sets keeping all sequences per subject while having different subjects in each set. In this way, there is no ground truth data to perform a quantitative evaluation for which EMOTE only conducted a perceptual user study. In contrast, our proposed split keeps a small number of sequences from each training subject for validation and testing allowing comparison of generated samples with respect to the ground truth. Although our training is done on fewer sequences than EMOTE, given the large scale of 3DMEAD dataset, we demonstrate that this split provides sufficient information for effective training and generation of animations perceptually superior in comparison to the EMOTE model. More details about the dataset split can be found in the supplementary material.
@@ 3.2. Problem Formulation
Figure 3. Stage 2: We keep the motion autoencoder trained in the previous stage frozen and train the HuBERT based Audio Encoder in such a way that given audio $a$, it produces the quantized audio latent, $z^{\prime\text{a}}$ from audio encoder output $z^{\text{a}}$ such that $z^{\prime\text{a}}$ closely resembles the quantized motion latent, $z^{\prime\text{m}}$. The Style Vector, $\mathcal{C}$ (i.e. concatenated one-hot vectors containing information about subject ID, emotion class and emotion intensity) is used to learn the Style Embedding that is fused with the encoded audio information. Unlike stage 1 training, the audio latent $z^{\text{a}}$ is used to get the quantized codebook latent $z^{\prime\text{a}}$ that is then decoded into facial animation $\hat{\mathcal{X}}$, utilizing the frozen motion decoder.
@@ 3.3.2. Stage 2: Speech and Emotion Conditioned
After learning the motion prior in stage 1, speech and emotion conditioned stage 2 is trained which consists of an audio encoder that encodes the continuous raw audio/speech data into discrete hidden representations and fuses the speaking style (i.e. style embedding obtained using Style Vector, $\mathcal{C}$) with the encoded audio information. As shown in Fig.3, the Motion Autoencoder trained in stage 1 is kept frozen in this stage to train the Audio Encoder in a supervised manner so that the latent representation, $z^{\text{a}}$ from the audio encoder produces a quantized latent $z^{\prime\text{a}}$ that closely resembles the quantized latent motion representation, $z^{\prime\text{m}}$.
@@
We employ pretrained HuBERT (Hsu et al., 2021), a transformer based speech recognition model in our audio encoder, and made appropriate modifications to apply it to our downstream task of facial animation synthesis following (Haque and Yumak, 2023; Stan et al., 2023). After the last hidden layer of HuBERT, inspired by the aforementioned previous works, we adjust the input representation so it temporally aligns with the paired animation data and add a linear projection layer that projects the encoded hidden state into a 256-dimensional hidden state to be element-wisely multiplied with the Style Embedding. Style Embedding is obtained using a linear projection of concatenated one-hot vectors of the style annotations (subject ID, emotion, intensity). In our case, the style vector, $\mathcal{C}$ is a 43-dimensional vector (as we have 32 subjects, 8 basic emotions, and 3 emotion intensity categories) that is linearly transformed into a 256-dimensional Style Embedding to facilitate the fusion of audio and style information. After fusing the audio information and style information, we introduce a 1D convolutional layer followed by 12 transformer layers to get the latent representation $z^{\text{a}}$ that goes through the prior codebook quantization and the motion decoder to obtain the predicted motion, $\hat{\mathcal{X}}$.
@@
The goal in stage 2 training is to get audio latent $z^{\text{a}}$ such that after quantization it produces $z^{\prime\text{a}}$ that closely resembles the quantized motion latent, $z^{\prime\text{m}}$. We construct $\mathcal{L}_{lat}=\mathcal{L}_{1}(z^{\prime\text{m}},z^{\prime\text{a}})$ that computes the $\mathcal{L}_{1}$ loss between these quantized latent representations. Additionally, we further add the reconstruction losses. The overall loss function is then defined in Eq.(5). The loss weights set for training are $\lambda_{\text{lat}}=1$, $\lambda_{\text{rec}}^{\text{exp}}=0.15$, and $\lambda_{\text{rec}}^{\text{jaw}}=0.1$.
@@ 4.1. Quantitative Evaluation
Figure 6. Comparison using heatmap visualization of mean and standard deviation of generated animation by different models together with ground truth (GT) given audio sequence, uttering the sentence: “He ate four extra eggs for breakfast”.